How much charge flows from each terminal of a 12.0-V battery when it is connected to a 5.00-μF capacitor?



1
Expert's answer
2022-03-10T11:15:59-0500

The charge is given as the product of capacitance and voltage:


"Q= CV = 5\times 10^{-6}F\cdot 12V = 2.160\times 10^{-2}C = 60.0\times 10^{-6}C"

Answer. "6.00\times 10^{-5}C".